OpenStudy (anonymous):

Algebra Help, please, I'm desperate! Given the function f(x) = 3x, find the value of f−1(81).

OpenStudy (anonymous):

f−1(81) = 1 f−1(81) = 2 f−1(81) = 3 f−1(81) = 4